40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 21 15:21:07 2013, yuta, Update, VAC, all heavy doors on, started pumping down annulus 
    Reply  Thu Feb 21 16:53:37 2013, Manasa, Update, PSL, PSL shutter 
       Reply  Fri Feb 22 13:03:49 2013, Manasa, Update, , MC REFL PD murder dead.jpg
          Reply  Fri Feb 22 17:33:25 2013, Manasa, Update, Electronics, MC REFL PD back from the dead 
             Reply  Wed Apr 22 17:04:11 2015, rana, Update, Electronics, MC REFL PD back from the dead 
                Reply  Thu Apr 23 08:43:40 2015, Steve, Update, Electronics, EG&G InGaAs diodes in stock 
    Reply  Thu Feb 21 18:01:27 2013, Steve, Update, VAC, pumpdown at 230 Torr 
       Reply  Thu Feb 21 21:12:42 2013, Koji, Update, VAC, pumpdown at 230 Torr 
          Reply  Fri Feb 22 08:01:25 2013, Steve, Update, VAC, pumpdown continuous 16hrs_pd75.png
             Reply  Fri Feb 22 11:12:16 2013, Steve, Update, VAC, pumpdown completed mag60min.pngpd75at20h.png
                Reply  Mon Feb 25 08:22:03 2013, Steve, Update, VAC, pumpdown at day 4 4d.png66d@atm.png
                   Reply  Fri Mar 1 14:50:36 2013, Steve, Update, VAC, pumpdown day 8 bgTP3d7.pngRGAonIFO.png
                      Reply  Wed Mar 6 08:40:03 2013, Steve, Update, VAC, first scan of this pumpdown pd75m12d.pngpd75mRGA12d.png
                         Reply  Mon Jun 3 09:46:40 2013, Steve, Update, VAC, RGA scan at day 100 d100Mpd75.pngd100Mpd75p.pngbadguys.png
                            Reply  Mon Jul 15 14:12:15 2013, Steve, Update, VAC, RGA scan at day 144 pd75mRGA144d.pngpd75m144d.png
Message ID: 8137     Entry time: Fri Feb 22 13:03:49 2013     In reply to: 8130     Reply to this: 8138
Author: Manasa 
Type: Update 
Category:  
Subject: MC REFL PD murder 

 

[Yuta, Manasa]

We turned IFO power back to 1.25W by removing the attenuator and forgot that the Y1 mirror before the REFL PD must be replaced with BS 10% before getting to full power. The PD is dead  and now we are in the process of fixing it. Forgive us for all our sins!

On the other note, we have changed the mech shutter mode from N.O back to N.C. So the shutter now works as usual from the medm screen.

dead.jpg

ELOG V3.1.3-